# ResQuNNs:Towards Enabling Deep Learning in Quantum Convolution Neural Networks

### v3

@@ III Experimental Setup
where $H(p,q)$ is the cross entropy loss, $p(c)$ represents the true probability distribution for class $c$ and $q(c)$ denotes the predicted probability distribution for class. All experiments are executed utilizing Pennylane, a Python library tailored for differentiable programming on quantum computing platforms [21]. An overview of the experimental toolflow utilized in this paper is depicted in Fig. 6.
@@
IV Results
IV-A Two Quanvolutional Layers
@@
IV-A1 Analysis of Gradients Accessibility
@@
IV-A2 Training Analysis
We then trained the proposed ResQuNNs on a subset of MNIST dataset, as detailed in Section III, for all the residual configurations labeled as 3 in Fig. 5, separately with classical and quantum postprocessing. The experimental procedures are consistent across all trials, adhering to the parameters discussed in Section III. It is worth mentioning here that the No residual configuration also serves as a common benchmark for all the other residual configurations to analyze whether or not the inclusion of residual connections bring any advantage. Below, we discuss the training results of ResQuNNs both with classical and quantum postprocessing.
@@ IV-A2 Training Analysis
  • No Residual. When there is No residual block, the QuNNs exhibit no training as shown in label 1 of Fig. 8.
  • X+O1. When the residual block is added after the first qunavolutional layer, i.e., the input is added to the output of first quanvolutional and then passed as input to the second quanvolutional layer, the model shows significant performance enhancement with improved training and validation accuracy compared to No residual setting, as shown in label 2 of Fig. 8. While $X+O1$ residual configuration marks a significant improvement over the No residual configuration, it nonetheless presents a sub-optimal training performance.
@@
Below, we scrutinize the influence of classical layer at the end in all these residual configurations to identify the residual configurations yielding better training and generalization performance where the quanvolutional layers primarily contribute to the learning.
@@
To assess the impact of the classical layer in various residual configurations of ResQuNNs, we conducted a detailed analysis using benchmark models. The benchmark models replicate the ResQuNNs with the residual configurations as shown in label 3 of Fig. 5, with one key difference: the quanvolutional layers were made untrainable, i.e., their weights are not updated during training, and only the classical layers were subjected to training. This exercise allows us to determine the extent to which the classical layer tends to dominate the learning process, compared to the quanvolutional layers across all configurations.
The benchmark models are trained using the same hyperparameters and dataset as detailed in Section III, and the results are compared with the models models, where both quanvolutional and classical layers were trainable. The comparison results are presented in Fig. 9. The observed performance trends in these models gives insights into the effectiveness of quanvolutional layers in different residual configurations. The key findings from this comparative analysis are summarized below:
@@
In this section, we present the training results of ResQuNNs using quantum layer/circuit to postprocess the results of quanvolutional layer(s), as shown in label 3 of Fig. 5. This very architectural configuration, we observe clear advantage of residual configurations that allow the gradient flow across both the layers of the network compared to those which allow partial gradient flow. The training results are depicted in Fig. 10. Below, we separately discuss the results of ResQuNN with different residual configurations (from label 5 in Fig. 5), employing two quanvolutional layers.
@@
IV-B Three Quanvolutional Layers
@@
Figure 11: Gradients Propagation Through the Network for Three Quanvolutional Layers for (O1+O2)+O3 (top) and ((X+O1)+O2)+O3 (bottom) and . To demonstrate gradients accessibility, we consider each quanvolutional layer with a quantum circuit depth of one
V Conclusion
In this paper, we presented a novel framework for developing Quantum Convolutional Neural Networks (QuNNs) by introducing trainable quanvolutional layers and the innovative concept of Residual Quanvolutional Neural Networks (ResQuNNs) for enabling the deep learning in QuNNs. Our approach addresses the significant challenges of gradient flow and adaptability in deep QuNN architectures, enabling enhanced training performance and scalability.

### v4

@@ IV Experimental Setup
where $H(p,q)$ is the cross entropy loss, $p(c)$ represents the true probability distribution for class $c$ and $q(c)$ denotes the predicted probability distribution for class. All experiments are executed utilizing Pennylane, a Python library tailored for differentiable programming on quantum computing platforms [27].
@@
V Results and Discussion
V-A Two Quanvolutional Layers
@@
V-A1 Analysis of Gradients Accessibility
@@
V-A2 Analysis of Training Performance
We then trained the proposed ResQuNNs on a subset of MNIST dataset, as detailed in Section IV, for all the residual configurations labeled as 3 in Fig. 5, separately with classical and quantum postprocessing. The experimental procedures are consistent across all trials, adhering to the parameters discussed in Section IV. It is worth mentioning here that the No residual configuration also serves as a common benchmark for all the other residual configurations to analyze whether or not the inclusion of residual connections bring any advantage. Below, we separately discuss the training results of ResQuNNs both with classical and quantum postprocessing.
@@ V-A2 Analysis of Training Performance
  • No Residual. When there is No residual block, the QuNNs exhibit no training as shown in label 1 of Fig. 8. Although, even in case of no residual, the gradients of second quanvolution layer are available but a significant information is lost in the first quanvolution layer (firstly because of the $2\times 2$ quanvolution filter, and secondly due to the measurements and re-encodings), which does not contribute to the learning because of no gradients.
  • X+O1. When the residual block is added after the first qunavolutional layer, i.e., the input is added to the output of first quanvolutional and then passed as input to the second quanvolutional layer, the model shows significant performance enhancement with improved training and validation accuracy compared to No residual setting, as shown in label 2 of Fig. 8. While $X+O1$ residual configuration marks a significant improvement over the No residual configuration, it nonetheless presents a sub-optimal training performance as the first quanvolutional layer does not take part in the overall optimization because of no gradients.
@@
Therefore, in next section, we scrutinize the influence of classical layer (used for postprocessing) on the overall learning performance in all the residual configurations from Fig. 5 label 5. This analysis is to identify the residual configurations yielding better training and generalization performance primarily due to the learning contributions from the quanvolutional layers.
@@
To assess the impact of the classical layer in various residual configurations of ResQuNNs, we conducted a detailed analysis using benchmark models. The benchmark models replicate the ResQuNNs with the residual configurations as shown in label 5 of Fig. 5, with one key difference: the quanvolutional layers were made untrainable, i.e., their weights are not updated during training, and only the classical layers were subjected to training. This exercise allows us to determine the extent to which the classical layer tends to dominate the learning process, compared to the quanvolutional layers across all the residual configurations. To put it simply, configurations yielding similar performance when the quanvolutional layer remains untrained and only the classical layer is trained owe their success primarily to the classical layer. Conversely, configurations that demonstrate improved performance when both the quanvolutional and classical layers are trained, as opposed to when only the classical layers are trained, benefit from the residual configuration and enhanced gradient flow throughout the network.
The benchmark models are trained using the same hyperparameters and dataset as detailed in Section IV, and the results are compared with the models, where both quanvolutional and classical layers were trainable. The comparison results are presented in Fig. 9. The observed performance trends in these models gives insights into the effectiveness of quanvolutional layers in different residual configurations. The key findings from this comparative analysis are summarized below:
@@
In this section, we present the training results of ResQuNNs using quantum layer/circuit to postprocess the results of quanvolutional layer(s), as shown in label 3 of Fig. 5. The primary reason behined quantum postprocessing is to reduce the influence of classical layer on the overall learning performance. While using quantum postprocessing, we observe clear advantage of residual configurations that allow the gradient flow across both the layers of the network compared to those which allow partial gradient flow. The training results are depicted in Fig. 10. Below, we separately discuss the results of ResQuNN with different residual configurations (from label 5 in Fig. 5), employing two quanvolutional layers.
@@
V-B Three Quanvolutional Layers
@@
Figure 11: Gradients Propagation Through the Network for Three Quanvolutional Layers for (O1+O2)+O3 (top) and ((X+O1)+O2)+O3 (bottom)
VI Conclusion
In this paper, we presented a novel framework for developing Quanvolutional Neural Networks (QuNNs) by introducing trainable quanvolutional layers and the concept of Residual Quanvolutional Neural Networks (ResQuNNs) for enabling the deep learning in QuNNs. Our approach addresses the significant challenges of gradient flow and adaptability in deep QuNN architectures, enabling enhanced training performance and scalability.